OpenStudy (anonymous):

How do I find the value of (-16)^1/4?

OpenStudy (anonymous):

You can google this stuff ya know, google is a calculator as well as many other things

OpenStudy (anonymous):

My answer is wrong. Btw. The answer is sqrt 2 + sqrt 2i My bad.

OpenStudy (anonymous):

its -4

OpenStudy (anonymous):

No, it's not. :) The fourth root of a negative number must contain i.

OpenStudy (anonymous):

its -2

OpenStudy (anonymous):

No it's not. It's sqrt 2 + sqrt 2i Google "Fourth root of -16" :)

OpenStudy (anonymous):

-16^0.25

OpenStudy (yanasidlinskiy):

http://www.wolframalpha.com/input/?i=%28-16%29%5E1%2F4%3F

OpenStudy (anonymous):

Stop googling for answers, if you don't know how to give them in proper syntax @VectorPrime :)

OpenStudy (anonymous):

I suck at math btw, dont trust my answers on here

OpenStudy (anonymous):

so whats the answer @ApoorvaAnand

OpenStudy (jdoe0001):

\(\Large { \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}}\qquad thus \\ \quad \\ (-16)^{\frac{{\color{blue}{ 1}}}{{\color{red}{ 4}}}}\implies \sqrt[{\color{red}{ 4}}]{-16^{\color{blue}{ 1}}} }\)

OpenStudy (jdoe0001):

any ideas on simplifying 16 to an value with an exponent?

OpenStudy (anonymous):

but wouldn't it be 16i

OpenStudy (jdoe0001):

well... sorta, not quite if you leave it at as 16... it'd be then .. one sec

OpenStudy (anonymous):

would it be 2 + sqrt(i)

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

I assume you don't have any choices?

OpenStudy (anonymous):

no i dont sorry

OpenStudy (jdoe0001):

I can make out this much thus far \(\bf \large { (-16)^{\frac{{\color{blue}{ 1}}}{{\color{red}{ 4}}}}\implies \sqrt[{\color{red}{ 4}}]{-16^{\color{blue}{ 1}}}\implies \sqrt[4]{-1\cdot 16}\implies \sqrt[4]{-1}\cdot \sqrt[4]{16} \\ \quad \\ \sqrt[4]{-1}\cdot \sqrt[4]{2^4}\implies 2\sqrt[4]{-1} }\)

OpenStudy (anonymous):

oh ok thnx, i got it cx

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

isnt 4radical(-1) the same as i^4

OpenStudy (anonymous):

@jdoe0001